# Near-optimal convergence of the full orthogonalization method

### v1

@@ Abstract.
2020 Mathematics Subject Classification:
Primary 65F10, 65F50
1. Introduction
The full orthogonalization method (FOM) [20] and the generalized minimal residual method (GMRES) [22] are two Krylov subspace methods used for solving a non-symmetric [^1] If $\mathbf{A}$ is symmetric GMRES is mathematically equivalent to MINRES [19], and if $\mathbf{A}$ is symmetric positive definite FOM is mathematically equivalent to conjugate gradient [14]. While this is relevant for efficient implementation, it does not impact the exact arithmetic theory in this note. linear system of equations
(1.1) $\mathbf{A}\mathbf{x}=\mathbf{b}.$
@@ 1. Introduction
(1.2) $\mathcal{K}_{k}(\mathbf{A},\mathbf{b}):=\operatorname{span}\{\mathbf{b},\mathbf{A}\mathbf{b},\ldots,\mathbf{A}^{k-1}\mathbf{b}\}.$
@@
(1.3) $\mathbf{A}\mathbf{Q}_{k}=\mathbf{Q}_{k+1}\mathbf{H}_{k+1,k}.$
@@
(1.4) $\mathbf{x}_{k}^{\mathrm{F}}:=\|\mathbf{b}\|_{2}\mathbf{Q}_{k}(\mathbf{H}_{k})^{-1}\mathbf{e}_{1},\qquad\mathbf{x}_{k}^{\mathrm{G}}:=\|\mathbf{b}\|_{2}\mathbf{Q}_{k}(\mathbf{H}_{k+1,k})^{\dagger}\mathbf{e}_{1},$
@@
(1.5) $\mathbf{r}_{k}^{\mathrm{F}}:=\mathbf{b}-\mathbf{A}\mathbf{x}_{k}^{\mathrm{F}},\qquad\mathbf{r}_{k}^{\mathrm{G}}:=\mathbf{b}-\mathbf{A}\mathbf{x}_{k}^{\mathrm{G}}.$
The norms of these residuals can be used as measure of how well the iterates $\mathbf{x}_{k}^{\mathrm{F}}$ and $\mathbf{x}_{k}^{\mathrm{G}}$ solve the linear system Equation 1.1, and understanding their relationship is the aim of this paper.
Figure 1. Residual norms for FOM $\|\mathbf{r}_{k}^{\mathrm{F}}\|_{2}$ (solid) and GMRES $\|\mathbf{r}_{k}^{\mathrm{G}}\|_{2}$ (dash-dot) for a symmetric matrix with eigenvalues in $[-10,-1]\cup[1,20]$ and the steam2 matrix. All curves are normalized by $\|\mathbf{b}\|_{2}$. While the residual norms for FOM jump up and down, they exhibit a generally downward trend which mirrors the convergence of the GMRES residual norms.
@@
(1.6) $\mathbf{x}_{k}^{\mathrm{G}}=\operatornamewithlimits{argmin}_{\mathbf{x}\in\mathcal{K}_{k}(\mathbf{A},\mathbf{b})}\|\mathbf{b}-\mathbf{A}\mathbf{x}\|_{2}.$
Hence, the GMRES residual norms are non-increasing and are optimal among Krylov subspace methods. This optimality guarantee leads to well-known results on the rate of convergence of the residual norms in terms of quantities such as the condition number of $\mathbf{A}$ [21]. On the other hand, the FOM residual norms often appear oscillatory, with large jumps. In fact, $\mathbf{H}_{k}$ may have an eigenvalue near or at zero, in which case $\|\mathbf{r}_{k}^{\mathrm{F}}\|_{2}$ can be arbitrarily large!
@@
2. Convergence bounds
Figure 2. Best FOM residual so far $\min_{j\leq k}\|\mathbf{r}_{j}^{\mathrm{F}}\|_{2}$ (solid), bound of Theorem 2.1 (dashed), and residual norms for FOM $\|\mathbf{r}_{k}^{\mathrm{F}}\|_{2}$ (solid grey) and GMRES $\|\mathbf{r}_{k}^{\mathrm{G}}\|_{2}$ (dash-dot). All curves are normalized by $\|\mathbf{b}\|_{2}$. While FOM exhibits oscillatory convergence, Theorem 2.1 ensures that the best FOM residual norm seen so far matches closely with the GMRES residual norm.
@@ 2. Convergence bounds
The theorem asserts that, while the FOM residual norm at any given iteration $k$ can be arbitrarily large, the overall convergence of FOM is at most $\sqrt{k+1}$ worse than that of GMRES. In light of the optimality of the GMRES iterates Equation 1.6, this implies the FOM iterates are, in a certain sense, near-optimal. While an immediate consequence of existing work, we have not been able to find this bound in the literature; see Subsection 3.1 for a discussion on past work. As we discuss in Section 3, Theorem 2.1 has implications for commonly used Krylov subspace methods for matrix functions.
In many practical situations, the convergence of GMRES is exponential in $k$, in which case the factor $\sqrt{k+1}$ is relatively small. In Figure 2, we show the examples from Figure 1 along with the bound from Theorem 2.1. To emphasize the “overall convergence” of FOM, we also display the smallest residual norm of FOM seen up to a given iteration $k$: $\min_{j\leq k}\|\mathbf{r}_{j}^{\mathrm{F}}\|_{2}$. As expected, this quantity tracks very closely the convergence of GMRES.
@@
(2.1) $\vartheta_{k+1}=-\frac{1}{h_{k+1,k}}\sum_{j=1}^{k}\vartheta_{j}h_{j,k},$
@@
Proof of Theorem 2.1.
Note that $\mathbf{H}_{k}$ is singular if and only if $|\vartheta_{k+1}|=0$. In this case, the FOM iterate is undefined and we can define the FOM residual norm error to be infinite. Therefore, defining $1/\infty=0$, we have that
(2.2) $\frac{\|\mathbf{r}_{k}^{\mathrm{G}}\|_{2}}{\|\mathbf{r}_{0}^{\mathrm{G}}\|_{2}}=\Bigg{(}\sum_{j=0}^{k}\frac{\|\mathbf{r}_{0}^{\mathrm{F}}\|_{2}^{2}}{\|\mathbf{r}_{j}^{\mathrm{F}}\|_{2}^{2}}\Bigg{)}^{-1/2}.$
@@
(2.3) $\|\mathbf{r}_{k}^{\mathrm{G}}\|_{2}=\Bigg{(}\sum_{j=0}^{k}\frac{1}{\|\mathbf{r}_{j}^{\mathrm{F}}\|_{2}^{2}}\Bigg{)}^{-1/2}.$
@@
(2.4) $\|\mathbf{r}_{k}^{\mathrm{G}}\|_{2}\geq\Bigg{(}(k+1)\max_{0\leq j\leq k}\frac{1}{\|\mathbf{r}_{j}^{\mathrm{F}}\|_{2}^{2}}\Bigg{)}^{-1/2}=\frac{1}{\sqrt{k+1}}\cdot\min_{0\leq j\leq k}\|\mathbf{r}_{j}^{\mathrm{F}}\|_{2},$
which proves the result. ∎
@@
For every $k\geq 1$ and $\varepsilon>0$, there exists a matrix $\mathbf{A}$ and vector $\mathbf{b}$ for which
$\min_{j\leq k}\|\mathbf{r}_{j}^{\mathrm{F}}\|_{2}\geq\left(\sqrt{k+1}-\varepsilon\right)\cdot\|\mathbf{r}_{k}^{\mathrm{G}}\|_{2}.$
Proof.
It is well-known that any sequence of residual norms is possible for FOM [18, §3.5]. In particular, fix $n>k$ and $\alpha\in(0,1)$. Then there exists a matrix $\mathbf{A}$ and vector $\mathbf{b}$ for which $\|\mathbf{r}_{j}^{\mathrm{F}}\|_{2}^{2}=\alpha^{j}$ for $j=0,1,\ldots,n$. Then $\|\mathbf{r}_{k}^{\mathrm{F}}\|_{2}=\min_{0\leq j\leq k}\|\mathbf{r}_{j}^{\mathrm{F}}\|_{2}$ and Equation 2.3 implies
$\|\mathbf{r}_{k}^{\mathrm{G}}\|_{2}^{2}=\frac{1}{\sum_{j=0}^{k}1/\|\mathbf{r}_{j}^{\mathrm{F}}\|_{2}^{2}}=\frac{1}{\sum_{j=0}^{k}\alpha^{-j}}=\frac{(\alpha-1)\alpha^{k}}{\alpha^{k+1}-1}=\frac{\alpha-1}{\alpha^{k+1}-1}\|\mathbf{r}_{k}^{\mathrm{F}}\|_{2}^{2}.$
Therefore, since $\lim_{\alpha\to 1}(\alpha-1)/(\alpha^{k+1}-1)=1/(k+1)$ and the above equation holds for any $\alpha\in(0,1)$, the result is proved. ∎
3. Discussion
3.1. Past work
@@
(3.1) $\|\mathbf{r}_{k}^{\mathrm{F}}\|_{2}=\frac{\|\mathbf{r}_{k}^{\mathrm{G}}\|_{2}}{\sqrt{1-\left(\|\mathbf{r}_{k}^{\mathrm{G}}\|_{2}/\|\mathbf{r}_{k-1}^{\mathrm{G}}\|_{2}\right)^{2}}},$
which can be obtained by rearranging Equation 2.3. Informally, Equation 3.1 says that at iterations where GMRES makes good progress (i.e. for which $\|\mathbf{r}_{k}^{\mathrm{G}}\|_{2}\ll\|\mathbf{r}_{k-1}^{\mathrm{G}}\|_{2}$) the FOM residual norm is close to the GMRES residual norm, and at iterations where GMRES stagnates (i.e. for which $\|\mathbf{r}_{k}^{\mathrm{G}}\|_{2}\approx\|\mathbf{r}_{k-1}^{\mathrm{G}}\|_{2}$) the FOM residual norm is very large. However, it does not provide any information about the iterations for which GMRES makes good progress, so it is unclear from Equation 3.1 that FOM’s overall convergence must track that of GMRES.
There are at least two bounds reminiscent of Theorem 2.1 in that they aim to understand the “overall convergence” of FOM, rather than the convergence at every iteration [13, 5]. While [13] uses entirely different techniques, [5] uses Equation 3.1 to argue that if GMRES is making good progress, there must be some iteration for which $\|\mathbf{r}_{j}^{\mathrm{G}}\|_{2}$ is substantially smaller than $\|\mathbf{r}_{j-1}^{\mathrm{G}}\|_{2}$ and hence for which $\|\mathbf{r}_{k}^{\mathrm{G}}\|_{2}\approx\|\mathbf{r}_{k-1}^{\mathrm{G}}\|_{2}$. Both bounds are weaker than Theorem 2.1 in that they (i) hold only for symmetric matrices and (ii) are not in terms of the GMRES residual norm, but rather the best approximation of zero on sets of the form $[a,b]\cup[c,b]$ by a polynomial taking value one at the origin, where $b<0<c$.
3.2. Error norm
@@
(3.2) $\min_{j\leq k}\|\mathbf{A}^{-1}\mathbf{b}-\mathbf{x}_{j}^{\mathrm{F}}\|_{2}\leq\sqrt{k+1}\cdot\kappa(\mathbf{A})\cdot\|\mathbf{A}^{-1}\mathbf{b}-\mathbf{x}_{k}^{\mathrm{G}}\|_{2}.$
@@
Figure 3. Best FOM error so far $\min_{j\leq k}\|\mathbf{A}^{-1}\mathbf{b}-\mathbf{x}_{k}^{\mathrm{F}}\|_{2}$ (solid), error for FOM $\|\mathbf{A}^{-1}\mathbf{b}-\mathbf{x}_{k}^{\mathrm{F}}\|_{2}$ (solid grey) and GMRES $\|\mathbf{A}^{-1}\mathbf{b}-\mathbf{x}_{k}^{\mathrm{G}}\|_{2}$ (dash-dot). All curves are normalized by $\|\mathbf{A}^{-1}\mathbf{b}\|_{2}$. Note that the FOM convergence is slightly better than GMRES, indicating FOM may be preferable to GMRES in some situations.
3.3. Connection to matrix functions
@@
(3.3) $f(x)=\int_{\Gamma}g(x)(x-z)^{-1}\mathrm{d}{x},$
for some weight function $g(x)$ and contour $\Gamma$. Common examples of functions which can be expressed in the form Equation 3.3 include analytic functions (exponential, indicator function for a region in the complex plane) for which $\Gamma$ is a contour in the complex plane, Stieltjes functions (inverse square root, logarithm) for which $\Gamma$ is a subset of the real line, and rational functions with simple poles, which are a special case of Stieltjes functions; see for instance [10]. Arnoldi-FA is widely use to compute the action of the corresponding matrix functions on a vector in applications throughout the computational sciences, including quantum chromodynamics, differential equations, machine learning, etc. [10, 15].
When $f(x)$ can be written as Equation 3.3 at the eigenvalues of $\mathbf{A}$ and $\mathbf{H}_{k}$, the Arnoldi-FA error is expressed as
(3.4) $f(\mathbf{A})\mathbf{b}-\mathbf{Q}_{k}f(\mathbf{H}_{k})\mathbf{e}_{1}=\int_{\Gamma}\Big{[}(\mathbf{A}-z\mathbf{I})^{-1}\mathbf{b}-\mathbf{Q}(\mathbf{H}_{k}-z\mathbf{I})^{-1}\mathbf{e}_{1}\Big{]}\,\mathrm{d}\mu(z).$

### v2

@@ Abstract
keywords:
Full Orthogonalization Method, GMRES
{AMS}
65F10
1 Introduction
The full orthogonalization method (FOM) [20] and the generalized minimal residual method (GMRES) [22] are two Krylov subspace methods used for solving a real or complex non-symmetric***If $\mathbf{A}$ is symmetric GMRES is mathematically equivalent to MINRES [19], and if $\mathbf{A}$ is symmetric positive definite FOM is mathematically equivalent to conjugate gradient [14]. While this is relevant for efficient implementation, it does not impact the exact arithmetic theory in this note. linear system of equations
(1) $\mathbf{A}\mathbf{x}=\mathbf{b}.$
@@ 1 Introduction
(2) $\mathcal{K}_{k}(\mathbf{A},\mathbf{b}):=\operatorname{span}\{\mathbf{b},\mathbf{A}\mathbf{b},\ldots,\mathbf{A}^{k-1}\mathbf{b}\}.$
@@
(3) $\mathbf{A}\mathbf{Q}_{k}=\mathbf{Q}_{k+1}\mathbf{H}_{k+1,k}.$
@@
(4) $\mathbf{x}_{k}^{\mathrm{F}}:=\|\mathbf{b}\|_{2}\mathbf{Q}_{k}(\mathbf{H}_{k})^{-1}\mathbf{e}_{1},\qquad\mathbf{x}_{k}^{\mathrm{G}}:=\|\mathbf{b}\|_{2}\mathbf{Q}_{k}(\mathbf{H}_{k+1,k})^{\dagger}\mathbf{e}_{1},$
@@
(5) $\mathbf{r}_{k}^{\mathrm{F}}:=\mathbf{b}-\mathbf{A}\mathbf{x}_{k}^{\mathrm{F}},\qquad\mathbf{r}_{k}^{\mathrm{G}}:=\mathbf{b}-\mathbf{A}\mathbf{x}_{k}^{\mathrm{G}}.$
The norms of these residuals can be used as measure of how well the iterates $\mathbf{x}_{k}^{\mathrm{F}}$ and $\mathbf{x}_{k}^{\mathrm{G}}$ solve the linear system Equation 1, and understanding their relationship is the aim of this paper.
Figure 1: Residual norms for FOM $\|\mathbf{r}_{k}^{\mathrm{F}}\|_{2}$ (solid) and GMRES $\|\mathbf{r}_{k}^{\mathrm{G}}\|_{2}$ (dash-dot) for a symmetric matrix with eigenvalues in $[-10,-1]\cup[1,20]$ and the steam2 matrix. All curves are normalized by $\|\mathbf{b}\|_{2}$. While the residual norms for FOM jump up and down, they exhibit a generally downward trend which mirrors the convergence of the GMRES residual norms.
@@
(6) $\mathbf{x}_{k}^{\mathrm{G}}=\operatornamewithlimits{argmin}_{\mathbf{x}\in\mathcal{K}_{k}(\mathbf{A},\mathbf{b})}\|\mathbf{b}-\mathbf{A}\mathbf{x}\|_{2}.$
Hence, the GMRES residual norms are non-increasing and are optimal among Krylov subspace methods. This optimality guarantee leads to well-known results on the rate of convergence of the residual norms in terms of quantities such as the condition number of $\mathbf{A}$ [21]. On the other hand, the FOM residual norms often appear oscillatory, with large jumps. In fact, it is easy to construct examples for which $\|\mathbf{r}_{k}^{\mathrm{F}}\|_{2}/\|\mathbf{r}_{0}^{\mathrm{F}}\|_{2}$ can be arbitrarily large [18]!
@@
2 Convergence bounds
Figure 2: Best FOM residual so far $\min_{j\leq k}\|\mathbf{r}_{j}^{\mathrm{F}}\|_{2}$ (solid), bound of Theorem 2.1 (dashed), and residual norms for FOM $\|\mathbf{r}_{k}^{\mathrm{F}}\|_{2}$ (solid grey) and GMRES $\|\mathbf{r}_{k}^{\mathrm{G}}\|_{2}$ (dash-dot). All curves are normalized by $\|\mathbf{b}\|_{2}$. While FOM exhibits oscillatory convergence, Theorem 2.1 ensures that the best FOM residual norm seen so far matches closely with the GMRES residual norm.
@@ 2 Convergence bounds
The theorem asserts that, while the FOM residual norm at any given iteration $k$ can be arbitrarily large, the overall convergence of FOM is at most $\sqrt{k+1}$ worse than that of GMRES. In light of the optimality of the GMRES iterates Equation 6, this implies the FOM iterates are, in a certain sense, near-optimal. While an immediate consequence of existing work, we have not been able to find this bound in the literature; see Subsection 3.1 for a discussion on past work. As we discuss in Section 3, Theorem 2.1 has implications for commonly used Krylov subspace methods for matrix functions.
In many practical situations, the GMRES residual norm is exponential in $k$, in which case the factor $\sqrt{k+1}$ is comparatively unimportant. In Figure 2, we show the examples from Figure 1 along with the bound from Theorem 2.1. To emphasize the “overall convergence” of FOM, we also display the smallest residual norm of FOM seen up to a given iteration $k$: $\min_{j\leq k}\|\mathbf{r}_{j}^{\mathrm{F}}\|_{2}$. As expected, this quantity tracks very closely the convergence of GMRES.
@@
(7) $\vartheta_{k+1}=-\frac{1}{h_{k+1,k}}\sum_{j=1}^{k}\vartheta_{j}h_{j,k},$
@@
Proof of Theorem 2.1. Note that $\mathbf{H}_{k}$ is singular if and only if $\vartheta_{k+1}=0$. In this case, the FOM iterate is undefined and we can define the FOM residual norm to be infinite. Therefore, defining $1/\infty=0$, we have that
(8) $\frac{\|\mathbf{r}_{k}^{\mathrm{G}}\|_{2}}{\|\mathbf{r}_{0}^{\mathrm{G}}\|_{2}}=\Bigg{(}\sum_{j=0}^{k}\frac{\|\mathbf{r}_{0}^{\mathrm{F}}\|_{2}^{2}}{\|\mathbf{r}_{j}^{\mathrm{F}}\|_{2}^{2}}\Bigg{)}^{-1/2}.$
@@
(9) $\|\mathbf{r}_{k}^{\mathrm{G}}\|_{2}=\Bigg{(}\sum_{j=0}^{k}\frac{1}{\|\mathbf{r}_{j}^{\mathrm{F}}\|_{2}^{2}}\Bigg{)}^{-1/2}.$
@@
(10) $\|\mathbf{r}_{k}^{\mathrm{G}}\|_{2}\geq\Bigg{(}(k+1)\max_{0\leq j\leq k}\frac{1}{\|\mathbf{r}_{j}^{\mathrm{F}}\|_{2}^{2}}\Bigg{)}^{-1/2}=\frac{1}{\sqrt{k+1}}\cdot\min_{0\leq j\leq k}\|\mathbf{r}_{j}^{\mathrm{F}}\|_{2},$
[…]
@@
For every $k\geq 1$ there exists a matrix $\mathbf{A}$ and vector $\mathbf{b}$ for which
$\min_{j\leq k}\|\mathbf{r}_{j}^{\mathrm{F}}\|_{2}=\sqrt{k+1}\cdot\|\mathbf{r}_{k}^{\mathrm{G}}\|_{2}.$
Proof 2.4.
It is well-known that any sequence of residual norms is possible for FOM; that is, given any sequence of non-zero positive numbers $f_{0},f_{1},\ldots,f_{k}$, there exists a matrix $\mathbf{A}$ and vector $\mathbf{b}$ (of dimension $n=k+1$) for which $\|\mathbf{r}_{j}^{\mathrm{F}}\|_{2}=f_{j}$ [18, Theorem 3.15]. In particular, there exists a matrix $\mathbf{A}$ and vector $\mathbf{b}$ for which $\|\mathbf{r}_{j}^{\mathrm{F}}\|_{2}^{2}=1$ for $j=0,1,\ldots,n$. In this case, Equation 9 implies
$\|\mathbf{r}_{k}^{\mathrm{G}}\|_{2}^{2}=\frac{1}{\sum_{j=0}^{k}1/\|\mathbf{r}_{j}^{\mathrm{F}}\|_{2}^{2}}=\frac{1}{k+1},$
which, since $\min_{0\leq j\leq k}\|\mathbf{r}_{j}^{\mathrm{F}}\|_{2}=1$, gives the result.
3 Discussion
3.1 Past work
@@
(11) $\|\mathbf{r}_{k}^{\mathrm{F}}\|_{2}=\frac{\|\mathbf{r}_{k}^{\mathrm{G}}\|_{2}}{\sqrt{1-\left(\|\mathbf{r}_{k}^{\mathrm{G}}\|_{2}/\|\mathbf{r}_{k-1}^{\mathrm{G}}\|_{2}\right)^{2}}},$
which can be obtained by rearranging Equation 9. Informally, Equation 11 says that at iterations where GMRES makes good progress (i.e. for which $\|\mathbf{r}_{k}^{\mathrm{G}}\|_{2}\ll\|\mathbf{r}_{k-1}^{\mathrm{G}}\|_{2}$) the FOM residual norm is close to the GMRES residual norm, and at iterations where GMRES stagnates (i.e. for which $\|\mathbf{r}_{k}^{\mathrm{G}}\|_{2}\approx\|\mathbf{r}_{k-1}^{\mathrm{G}}\|_{2}$) the FOM residual norm is very large. However, it does not provide any information about at which, if any, iterations GMRES makes good progress. As such, it is unclear from Equation 11 that FOM’s overall convergence must track that of GMRES.
There are at least two bounds reminiscent of Theorem 2.1 in that they aim to understand the “overall convergence” of FOM, rather than the convergence at every iteration [13, 5]. While [13] uses entirely different techniques, [5] uses Equation 11 to argue that if GMRES is making good progress, there must be some iteration for which $\|\mathbf{r}_{j}^{\mathrm{G}}\|_{2}$ is substantially smaller than $\|\mathbf{r}_{j-1}^{\mathrm{G}}\|_{2}$ and hence for which $\|\mathbf{r}_{k}^{\mathrm{G}}\|_{2}\approx\|\mathbf{r}_{k-1}^{\mathrm{G}}\|_{2}$. Both bounds are weaker than Theorem 2.1 in that they (i) hold only for symmetric matrices and (ii) are not in terms of the GMRES residual norm, but rather the best approximation of zero on sets of the form $[a,b]\cup[c,b]$ by a polynomial taking value one at the origin, where $b<0<c$.
3.2 Error norm
@@
(12) $\min_{j\leq k}\|\mathbf{A}^{-1}\mathbf{b}-\mathbf{x}_{j}^{\mathrm{F}}\|_{2}\leq\sqrt{k+1}\cdot\kappa(\mathbf{A})\cdot\|\mathbf{A}^{-1}\mathbf{b}-\mathbf{x}_{k}^{\mathrm{G}}\|_{2}.$
@@
Figure 3: Best FOM error so far $\min_{j\leq k}\|\mathbf{A}^{-1}\mathbf{b}-\mathbf{x}_{k}^{\mathrm{F}}\|_{2}$ (solid), error for FOM $\|\mathbf{A}^{-1}\mathbf{b}-\mathbf{x}_{k}^{\mathrm{F}}\|_{2}$ (solid grey) and GMRES $\|\mathbf{A}^{-1}\mathbf{b}-\mathbf{x}_{k}^{\mathrm{G}}\|_{2}$ (dash-dot). All curves are normalized by $\|\mathbf{A}^{-1}\mathbf{b}\|_{2}$. Note that the FOM convergence is slightly better than GMRES, indicating FOM may be preferable to GMRES in some situations.
3.3 Connection to matrix functions
@@
(13) $f(x)=\int_{\Gamma}g(x)(x-z)^{-1}\mathrm{d}{x},$
for some weight function $g(x)$ and contour $\Gamma$. Common examples of functions which can be expressed in the form Equation 13 include analytic functions (exponential, indicator function for a region in the complex plane) for which $\Gamma$ is a contour in the complex plane, Stieltjes functions (inverse square root, logarithm) for which $\Gamma$ is a subset of the real line, and rational functions with simple poles, which are a special case of Stieltjes functions; see for instance [10]. Arnoldi-FA is widely used to compute the action of the corresponding matrix functions on a vector in applications throughout the computational sciences, including quantum chromodynamics, differential equations, machine learning, etc. [10, 15].
When $f(x)$ can be written as Equation 13 at the eigenvalues of $\mathbf{A}$ and $\mathbf{H}_{k}$, the Arnoldi-FA error is expressed as
(14) $f(\mathbf{A})\mathbf{b}-\mathbf{Q}_{k}f(\mathbf{H}_{k})\mathbf{e}_{1}=\int_{\Gamma}\Big{[}(\mathbf{A}-z\mathbf{I})^{-1}\mathbf{b}-\mathbf{Q}(\mathbf{H}_{k}-z\mathbf{I})^{-1}\mathbf{e}_{1}\Big{]}\,\mathrm{d}\mu(z).$